16. Use incremental analysis to select the best alternatives using Benefit-Cost ration 4000 2000 6000 7330 470008730 GO0D 1000 9000 10000 Cost Pw(benefits) 90009500 Useful life is 20years Interest 6%
Answer #1

Total opportunity cost of investment A = 4000(1 + 0.06)20 = 3.2071 × 4000 = 12828

Net benefit = 7330 - 12828 = -5498

Total opportunity cost of investment B = 2000(1 + 0.06)20= 6414.2

Net benefit = 47000 - 6414.2 = 40585.8

Total opportunity cost of investment C = 6000 (1 + 0.06)20= 19242.6

Net benefit = 8730 - 19242.6 = -10512.6

Total opportunity cost of investment D = 1000(1 +0.06)20 = 3207.1

Net benefit = 1340 - 3207.1 = - 1867.1

Total opportunity cost of investment E = 9000(1+0.06)20= 28863.9

Net benefit = 9000- 28863.9 = -19863

Total opportunity cost of investment F = 10000(1+0.06)20 = 32071

Net benefit = 9500 - 32071 = - 22571

Since the net benefit is negative in all investments except B, B is the best alternative.